# Distributionally Robust Beamforming and Estimation of Wireless Signals

### v1

@@ IV-A General Framework and Concrete Examples
We consider a distributionally robust Capon beamforming problem:
$\begin{array}[]{cl}\displaystyle\min_{\bm{W}}\max_{\bm{R}_{x}}&\operatorname{Tr}\big{[}\bm{W}\bm{R}_{x}\bm{W}^{\mathsf{H}}-2\bm{P}+\bm{R}_{s}\big{]}\\ \text{s.t.}&\bm{W}\bm{H}=\bm{I}_{M},\\ &{\hat{\bm{R}}}_{x}-\epsilon_{1}\bm{I}_{N}\preceq\bm{R}_{x}\preceq{\hat{\bm{R}}}_{x}+\epsilon_{1}\bm{I}_{N},\\ &\bm{R}_{x}\succeq\bm{0},~{}\bm{R}_{x}=\bm{R}^{\mathsf{H}}_{x},\end{array}$
@@
$\begin{array}[]{cl}\displaystyle\min_{\bm{W}}\max_{\bm{R}_{x}}&\operatorname{Tr}\big{[}\bm{W}\bm{R}_{x}\bm{W}^{\mathsf{H}}\big{]}\\ \text{s.t.}&\bm{W}\bm{H}=\bm{I}_{M},\\ &{\hat{\bm{R}}}_{x}-\epsilon_{1}\bm{I}_{N}\preceq\bm{R}_{x}\preceq{\hat{\bm{R}}}_{x}+\epsilon_{1}\bm{I}_{N},\\ &\bm{R}_{x}\succeq\bm{0},~{}\bm{R}_{x}=\bm{R}^{\mathsf{H}}_{x}.\end{array}$
According to Lemma 1 and Theorem 1, the above display is further equivalent to
$\begin{array}[]{cl}\displaystyle\min_{\bm{W}}&\operatorname{Tr}\big{[}\bm{W}({\hat{\bm{R}}}_{x}+\epsilon_{1}\bm{I}_{N})\bm{W}^{\mathsf{H}}\big{]}\\ \text{s.t.}&\bm{W}\bm{H}=\bm{I}_{M},\end{array}$
that is,
$\begin{array}[]{cl}\displaystyle\min_{\bm{W}}&\operatorname{Tr}\big{[}\bm{W}{\hat{\bm{R}}}_{x}\bm{W}^{\mathsf{H}}\big{]}+\epsilon_{1}\cdot\operatorname{Tr}\big{[}\bm{W}\bm{W}^{\mathsf{H}}\big{]}\\ \text{s.t.}&\bm{W}\bm{H}=\bm{I}_{M}.\end{array}$
@@
$\begin{array}[]{l}\bm{W}^{\star}_{\text{DR-Capon}}=[\bm{H}^{\mathsf{H}}(\hat{\bm{R}}_{x}+\epsilon_{1}\bm{I}_{N})^{-1}\bm{H}]^{-1}\cdot\\ \quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\bm{H}^{\mathsf{H}}(\hat{\bm{R}}_{x}+\epsilon_{1}\bm{I}_{N})^{-1},\end{array}$ (28)
which is the diagonally loaded Capon beamformer. normal-$\square$
@@
[…]
@@
[…]
which is known as the eigenvalue thresholding method [20], [14, Eq. (12)]. normal-$\square$
In the next example, we consider the uncertainties in covariance $\bm{P}$ of transmitted signal $\mathbf{s}$ and covariance $\bm{R}$ of channel noise $\mathbf{v}$.
Example 4 (Distributionally Robust Beamforming for Uncertain $\bm{P}$ and $\bm{R}$)
Consider Problem (24), as well as the uncertainties in covariance $\bm{P}$ of transmitted signal $\mathbf{s}$ and covariance $\bm{R}$ of channel noise $\mathbf{v}$. Since the objective function of (24) is increasing in both $\bm{P}$ and $\bm{R}$, [^1] This claim can be routinely proven in analogy to Theorem 1 and a real-space case in [29, Thm. 1]. if
$\hat{\bm{P}}-\epsilon_{4}\bm{I}_{M}\preceq\bm{P}\preceq\hat{\bm{P}}+\epsilon_{4}\bm{I}_{M},$
@@
$\begin{array}[]{cl}\bm{W}^{\star}_{\text{DR-P}}&=(\hat{\bm{P}}+\epsilon_{4}\bm{I}_{M})\bm{H}^{\mathsf{H}}[\bm{H}(\hat{\bm{P}}+\epsilon_{4}\bm{I}_{M})\bm{H}^{\mathsf{H}}+\bm{R}]^{-1}\\ &=(\hat{\bm{P}}+\epsilon_{4}\bm{I}_{M})\bm{H}^{\mathsf{H}}[\bm{H}\hat{\bm{P}}\bm{H}^{\mathsf{H}}+\bm{R}+\epsilon_{4}\bm{H}\bm{H}^{\mathsf{H}}]^{-1},\end{array}$ (31)
which is tantamount to the diagonally loaded beamformer but loaded by the squared channel matrix $\bm{H}\bm{H}^{\mathsf{H}}$ rather than the identity matrix. On the other hand, if
$\hat{\bm{R}}-\epsilon_{5}\bm{I}_{N}\preceq\bm{R}\preceq\hat{\bm{R}}+\epsilon_{5}\bm{I}_{N},$
we have a distributionally robust beamformer
$\bm{W}^{\star}_{\text{DR-R}}=\bm{P}\bm{H}^{\mathsf{H}}[\bm{H}\bm{P}\bm{H}^{\mathsf{H}}+\hat{\bm{R}}+\epsilon_{5}\bm{I}_{N}]^{-1},$ (32)
which is also a diagonally loaded beamformer. normal-$\square$
Example 5 (Generalized Diagonal Loading)
The generalized diagonally loaded beamformer can be obtained if $\bm{R}_{x}\preceq{\hat{\bm{R}}}_{x}+\epsilon\bm{F}$ or $\bm{R}\preceq{\hat{\bm{R}}}+\epsilon\bm{F}$ for some $\bm{F}\succeq\bm{0}$:
$\begin{array}[]{cl}\bm{W}^{\star}_{\text{DR-GDL}}&=\bm{R}^{\mathsf{H}}_{xs}[{\hat{\bm{R}}}_{x}+\epsilon\bm{F}]^{-1}\\ &=\bm{P}\bm{H}^{\mathsf{H}}[\bm{H}\bm{P}\bm{H}^{\mathsf{H}}+\hat{\bm{R}}+\epsilon\bm{F}]^{-1}.\end{array}$ (33)
When $\bm{H}$ is known, we can choose $\bm{F}\coloneqq\bm{H}\bm{H}^{\mathsf{H}}$, which is reminiscent of (31). normal-$\square$
From Examples 1-5, we can see that the popular diagonally loaded beamformer (27) and the proposed generalized diagonally loaded beamformer (33) can combat several types of uncertainties including the limited training sample size, the uncertainty in the covariance of the transmitted signal, the uncertainty in the covariance of the channel noise, etc. However, the uncertainty sets are quite specially constructed, e.g., using the diagonal $\epsilon$-perturbation method, the eigenvalue thresholding method, or the $\bm{F}$ matrix.
In addition, Example 2 motivates the following important theorem, which justifies the popular ridge regression in machine learning.
@@
${\hat{\bm{R}}}_{x}-\epsilon\bm{I}_{N}\preceq\bm{R}_{x}\preceq{\hat{\bm{R}}}_{x}+\epsilon\bm{I}_{N},$
@@
$\min_{\bm{W}}\operatorname{Tr}\big{[}\bm{W}\hat{\bm{R}}_{x}\bm{W}^{\mathsf{H}}-\bm{W}\hat{\bm{R}}_{xs}-\hat{\bm{R}}^{\mathsf{H}}_{xs}\bm{W}^{\mathsf{H}}+\hat{\bm{R}}_{s}\big{]}+\epsilon\operatorname{Tr}\big{[}\bm{W}\bm{W}^{\mathsf{H}}\big{]}.$
@@
${\hat{\bm{R}}}_{x}-\epsilon\bm{F}\preceq\bm{R}_{x}\preceq{\hat{\bm{R}}}_{x}+\epsilon\bm{F}$
is employed for some $\bm{F}\succeq\bm{0}$.
@@
This is due to Lemma 1 and Theorem 1. Just note that $\operatorname{Tr}\big{[}\bm{W}(\hat{\bm{R}}_{x}+\epsilon\bm{F})\bm{W}^{\mathsf{H}}-\bm{W}\hat{\bm{R}}_{xs}-\hat{\bm{R}}^{\mathsf{H}}_{xs}\bm{W}^{\mathsf{H}}+\hat{\bm{R}}_{s}\big{]}=\\ \operatorname{Tr}\big{[}\bm{W}\hat{\bm{R}}_{x}\bm{W}^{\mathsf{H}}-\bm{W}\hat{\bm{R}}_{xs}-\hat{\bm{R}}^{\mathsf{H}}_{xs}\bm{W}^{\mathsf{H}}+\hat{\bm{R}}_{s}\big{]}+\epsilon\operatorname{Tr}\big{[}\bm{W}\bm{F}\bm{W}^{\mathsf{H}}\big{]}$. This completes the proof. $\square$
Note that in Theorem 2, the second-order moment of $\mathbf{s}$ is not considered because it does not influence the optimal solution of $\bm{W}$: i.e., the optimal solution of $\bm{W}$ does not depend on the value of $\bm{R}_{s}$. Theorem 2 gives a theoretical interpretation of the popular ridge regression in machine learning from the perspective of distributional robustness. [^2] This is the first time to provide the ridge regression with an interpretation of distributional robustness. When the uncertainty is quantified by the Wasserstein distance, a similar result can be seen in [30, Prop. 3], which however is not a ridge regression formulation because in [30, Prop. 3], the loss function is square-rooted and the norm regularizer is not squared; cf. also [27, Rem. 18 and 19]. The corollary below justifies the rationale of any norm-regularized method.
@@
Motivated by Theorem 2, the following corollary is immediate, which gives another interpretation of ridge regression and Tikhonov regularization from the perspective of data augmentation through data perturbation (i.e., the noise injection method in machine learning).
@@
Suppose that $\mathbf{\Delta}_{x}$ is uncorrelated with $\mathbf{x}$, with $\mathbf{s}$, and with $\mathbf{\Delta}_{s}$; in addition, $\mathbf{\Delta}_{s}$ is uncorrelated with $\mathbf{x}$. If the second-order moment of $\mathbf{\Delta}_{x}$ is upper bounded as $\mathbb{E}\mathbf{\Delta}_{x}\mathbf{\Delta}^{\mathsf{H}}_{x}\preceq\epsilon\bm{I}_{N}$, then the distributionally robust linear estimator of $\bm{W}$ becomes a ridge regression (i.e., squared-$F$-norm regularized) method
$\min_{\bm{W}}\operatorname{Tr}\big{[}\bm{W}\hat{\bm{R}}_{x}\bm{W}^{\mathsf{H}}-\bm{W}\hat{\bm{R}}_{xs}-\hat{\bm{R}}^{\mathsf{H}}_{xs}\bm{W}^{\mathsf{H}}+\hat{\bm{R}}_{s}\big{]}+\epsilon\operatorname{Tr}\big{[}\bm{W}\bm{W}^{\mathsf{H}}\big{]}.$
The regularization term becomes $\operatorname{Tr}\big{[}\bm{W}\bm{F}\bm{W}^{\mathsf{H}}\big{]}$, which is known as the Tikhonov regularizer, if $\mathbb{E}\mathbf{\Delta}_{x}\mathbf{\Delta}^{\mathsf{H}}_{x}\preceq\epsilon\bm{F}$, for some $\bm{F}\succeq\bm{0}$. normal-$\square$
@@
IV-B Non-Trivial Uncertainty Sets
Below we study more general and non-trivial construction methods for the uncertainty sets of $\bm{R}_{x}$, $\bm{P}$, and $\bm{R}$ using the Wasserstein distance and the $F$-norm, beyond the simple diagonal $\epsilon$-perturbation method and the eigenvalue thresholding method.
@@ IV-B1 Wasserstein Distributionally Robust Beamforming
We start with the case of $\bm{R}_{x}$:
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}_{x}}&\operatorname{Tr}\big{[}-\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{-1}_{x}\bm{R}_{xs}+\bm{R}_{s}\big{]}\\ \text{s.t.}&\operatorname{Tr}\left[{\bm{R}_{x}+\hat{\bm{R}}_{x}-2(\hat{\bm{R}}^{1/2}_{x}\bm{R}_{x}\hat{\bm{R}}^{1/2}_{x})^{1/2}}\right]\leq\epsilon^{2}_{1}\\ &\bm{R}_{x}\succeq\bm{0},~{}\bm{R}_{x}=\bm{R}^{\mathsf{H}}_{x},\end{array}$ (35)
where we assumed that the pseudo-covariance (i.e., relation matrix) of the complex vector $\mathbf{x}$ has not been perturbed. Otherwise, supposing the pseudo-covariance of $\mathbf{x}$ is $\bm{C}_{x}$, we can consider one more constraint
$\operatorname{Tr}[\bm{C}_{x}+\hat{\bm{C}}_{x}-2(\hat{\bm{C}}^{1/2}_{x}\bm{C}_{x}\hat{\bm{C}}^{1/2}_{x})^{1/2}]\leq\epsilon^{2}_{1,2},$
where $\epsilon_{1,2}\geq 0$ is an uncertainty quantification parameter for the uncertainty in $\hat{\bm{C}}_{x}$.
@@
Problem (35) is equivalent to a linear P-SDP
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}_{x},\bm{V},\bm{U}}&\operatorname{Tr}\big{[}-\bm{V}+\bm{R}_{s}\big{]}\\ \text{s.t.}&\left[\begin{array}[]{cc}\bm{V}&\bm{R}^{\mathsf{H}}_{xs}\\ \bm{R}_{xs}&\bm{R}_{x}\end{array}\right]\succeq\bm{0}\\ &\operatorname{Tr}\left[{\bm{R}_{x}+\hat{\bm{R}}_{x}-2\bm{U}}\right]\leq\epsilon^{2}_{1}\\ &\left[\begin{array}[]{cc}\hat{\bm{R}}^{1/2}_{x}\bm{R}_{x}\hat{\bm{R}}^{1/2}_{x}&\bm{U}\\ \bm{U}&\bm{I}_{N}\end{array}\right]\succeq\bm{0}\\ &\bm{R}_{x}=\bm{R}^{\mathsf{H}}_{x},~{}\bm{V}=\bm{V}^{\mathsf{H}},~{}\bm{U}=\bm{U}^{\mathsf{H}}\\ &\bm{R}_{x}\succeq\bm{0},~{}\bm{V}\succeq\bm{0},~{}\bm{U}\succeq\bm{0}.\end{array}$ (36)
@@
See Appendix F of the online supplementary materials. $\square$
Complex-valued linear P-SDP can be solved using the YALMIP solver. [^3] See https://yalmip.github.io/inside/complexproblems/.
Suppose that $\bm{R}^{\star}_{x}$ solves (36). The corresponding Wasserstein distributionally robust beamformer is given as
$\bm{W}^{\star}_{\text{DR-Wasserstein}}=\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{*-1}_{x}.$ (37)
Next, we investigate the case of uncertain $\hat{\bm{P}}$ and $\hat{\bm{R}}$. From (24), we have
$\begin{array}[]{cl}\displaystyle\max_{\bm{P},\bm{R}}&\operatorname{Tr}\big{[}\bm{P}-\bm{P}\bm{H}^{\mathsf{H}}(\bm{H}\bm{P}\bm{H}^{\mathsf{H}}+\bm{R})^{-1}\bm{H}\bm{P}\big{]}\\ \text{s.t.}&\operatorname{Tr}\left[{\bm{P}+\hat{\bm{P}}-2(\hat{\bm{P}}^{1/2}\bm{P}\hat{\bm{P}}^{1/2})^{1/2}}\right]\leq\epsilon^{2}_{4}\\ &\operatorname{Tr}\left[{\bm{R}+\hat{\bm{R}}-2(\hat{\bm{R}}^{1/2}\bm{R}\hat{\bm{R}}^{1/2})^{1/2}}\right]\leq\epsilon^{2}_{5}\\ &\bm{P}=\bm{P}^{\mathsf{H}},~{}\bm{R}=\bm{R}^{\mathsf{H}}\\ &\bm{P}\succeq\bm{0},~{}\bm{R}\succeq\bm{0},\end{array}$ (38)
where we have assumed that the pseudo-covariances of the complex vectors $\mathbf{s}$ and $\mathbf{v}$ have not been perturbed.
Problem (38) can be transformed into a linear P-SDP using a similar technique in the proof of Proposition 1. One can just introduce an inequality $\bm{U}\succeq\bm{P}\bm{H}^{\mathsf{H}}(\bm{H}\bm{P}\bm{H}^{\mathsf{H}}+\bm{R})^{-1}\bm{H}\bm{P}$ and the objective function will become $\operatorname{Tr}\left[{\bm{P}-\bm{U}}\right]$.
Suppose that $(\bm{P}^{\star},\bm{R}^{\star})$ solves (38). The corresponding Wasserstein distributionally robust beamformer is given as
$\bm{W}^{\star}_{\text{DR-Wasserstein-PR}}=\bm{P}^{\star}\bm{H}^{\mathsf{H}}[\bm{H}\bm{P}^{\star}\bm{H}^{\mathsf{H}}+\bm{R}^{\star}]^{-1}.$ (39)
The Wasserstein distance is valid for Hermitian positive semi-definite matrices that are covariance matrices of Gaussian distributions. Therefore, we cannot use the Wasserstein distance to measure the differences between $\bm{R}_{xs}$ and $\hat{\bm{R}}_{xs}$, and between $\bm{H}$ and $\hat{\bm{H}}$. In this case, the $F$-norm of matrices can be used.

### v2

@@ 4.  C4)
We consider a distributionally robust Capon beamforming problem under the trimmed uncertainty set (34):
$\begin{array}[]{cl}\displaystyle\min_{\bm{W}}\max_{\bm{R}_{x}}&\operatorname{Tr}\big{[}\bm{W}\bm{R}_{x}\bm{W}^{\mathsf{H}}-2\bm{R}_{s}+\bm{R}_{s}\big{]}\\ \text{s.t.}&\bm{W}\bm{H}=\bm{I}_{M},\\ &{\hat{\bm{R}}}_{x}-\epsilon_{0}\bm{I}_{N}\preceq\bm{R}_{x}\preceq{\hat{\bm{R}}}_{x}+\epsilon_{0}\bm{I}_{N},\\ &\bm{R}_{x}\succeq\bm{0},\end{array}$
@@
$\begin{array}[]{cl}\displaystyle\min_{\bm{W}}\max_{\bm{R}_{x}}&\operatorname{Tr}\big{[}\bm{W}\bm{R}_{x}\bm{W}^{\mathsf{H}}\big{]}\\ \text{s.t.}&\bm{W}\bm{H}=\bm{I}_{M},\\ &{\hat{\bm{R}}}_{x}-\epsilon_{0}\bm{I}_{N}\preceq\bm{R}_{x}\preceq{\hat{\bm{R}}}_{x}+\epsilon_{0}\bm{I}_{N},\\ &\bm{R}_{x}\succeq\bm{0}.\end{array}$
According to Theorem 1, the above display is equivalent to
$\begin{array}[]{cl}\displaystyle\min_{\bm{W}}&\operatorname{Tr}\big{[}\bm{W}{\hat{\bm{R}}}_{x}\bm{W}^{\mathsf{H}}\big{]}+\epsilon_{0}\cdot\operatorname{Tr}\big{[}\bm{W}\bm{W}^{\mathsf{H}}\big{]}\\ \text{s.t.}&\bm{W}\bm{H}=\bm{I}_{M}.\end{array}$
@@
$\begin{array}[]{l}\bm{W}^{\star}_{\text{DR-Capon}}=[\bm{H}^{\mathsf{H}}(\hat{\bm{R}}_{x}+\epsilon_{0}\bm{I}_{N})^{-1}\bm{H}]^{-1}\cdot\\ \quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\bm{H}^{\mathsf{H}}(\hat{\bm{R}}_{x}+\epsilon_{0}\bm{I}_{N})^{-1},\end{array}$ (38)
which is the diagonal-loading Capon beamformer. $\square$
@@
Suppose that $\hat{\bm{R}}_{x}$ admits the eigenvalues of $\operatorname{diag}\{\lambda_{1},\lambda_{2},\ldots,\lambda_{N}\}$ in descending order and the eigenvectors in $\bm{Q}$ (columns). Let $0\leq\mu\leq 1$ be a shrinking coefficient. If we assume $\bm{R}_{x}\preceq\hat{\bm{R}}_{x,\text{thr}}$ where
$\hat{\bm{R}}_{x,\text{thr}}\coloneqq\bm{Q}{\footnotesize\left[\begin{array}[]{cccc}\lambda_{1}&&&\\ &\max\{\mu\lambda_{1},\lambda_{2}\}&&\\ &&\ddots&\\ &&&\max\{\mu\lambda_{1},\lambda_{N}\}\end{array}\right]}\bm{Q}^{-1},$ (39)
@@
$\bm{W}^{\star}_{\text{DR-ET}}=\bm{R}^{\mathsf{H}}_{xs}\hat{\bm{R}}^{-1}_{x,\text{thr}},$ (40)
which is known as the eigenvalue thresholding method [20], [14, Eq. (12)]. $\square$
Example 3 (Distributionally Robust Beamforming for Uncertain $\bm{R}_{s}$ and $\bm{R}_{v}$)
Consider Problem (36). Since the objective of (36) is increasing in both $\bm{R}_{s}$ and $\bm{R}_{v}$, [^1] This claim can be routinely proven in analogy to Theorem 1 and a real-space case in [30, Thm. 1]. if
$\hat{\bm{R}}_{s}-\epsilon_{1}\bm{I}_{M}\preceq\bm{R}_{s}\preceq\hat{\bm{R}}_{s}+\epsilon_{1}\bm{I}_{M},$
@@
$\begin{array}[]{cl}\bm{W}^{\star}_{\text{DR}}&=(\hat{\bm{R}}_{s}+\epsilon_{1}\bm{I}_{M})\bm{H}^{\mathsf{H}}[\bm{H}(\hat{\bm{R}}_{s}+\epsilon_{1}\bm{I}_{M})\bm{H}^{\mathsf{H}}+\bm{R}_{v}]^{-1}\\ &=(\hat{\bm{R}}_{s}+\epsilon_{1}\bm{I}_{M})\bm{H}^{\mathsf{H}}[\bm{H}\hat{\bm{R}}_{s}\bm{H}^{\mathsf{H}}+\bm{R}_{v}+\epsilon_{1}\bm{H}\bm{H}^{\mathsf{H}}]^{-1};\end{array}$ (41)
if instead
$\hat{\bm{R}}_{s}-\epsilon_{1}\bm{H}^{\mathsf{H}}(\bm{H}\bm{H}^{\mathsf{H}})^{-2}\bm{H}\preceq\bm{R}_{s}\preceq\hat{\bm{R}}_{s}+\epsilon_{1}\bm{H}^{\mathsf{H}}(\bm{H}\bm{H}^{\mathsf{H}})^{-2}\bm{H},$ (42)
we have
$\begin{array}[]{l}\bm{W}^{\star}_{\text{DR}}=[\hat{\bm{R}}_{s}\bm{H}^{\mathsf{H}}+\epsilon_{1}\bm{H}^{\mathsf{H}}(\bm{H}\bm{H}^{\mathsf{H}})^{-1}]\cdot\\ \quad\quad\quad\quad\quad[\bm{H}\hat{\bm{R}}_{s}\bm{H}^{\mathsf{H}}+\bm{R}_{v}+\epsilon_{1}\bm{I}_{M}]^{-1},\end{array}$ (43)
which is a modified diagonal-loading beamformer. On the other hand, if
$\hat{\bm{R}}_{v}-\epsilon_{2}\bm{I}_{N}\preceq\bm{R}_{v}\preceq\hat{\bm{R}}_{v}+\epsilon_{2}\bm{I}_{N},$
we have
$\bm{W}^{\star}_{\text{DR}}=\bm{R}_{s}\bm{H}^{\mathsf{H}}[\bm{H}\bm{R}_{s}\bm{H}^{\mathsf{H}}+\hat{\bm{R}}_{v}+\epsilon_{2}\bm{I}_{N}]^{-1},$ (44)
which is also a diagonal-loading beamformer. $\square$
Motivated by Corollary 1 and Examples 1$\sim$3, as well as the trimmed uncertainty sets in Definition 5, we have the following important theorem, which justifies the popular ridge regression in machine learning.
@@
${\hat{\bm{R}}}_{x}-\epsilon_{0}\bm{I}_{N}\preceq\bm{R}_{x}\preceq{\hat{\bm{R}}}_{x}+\epsilon_{0}\bm{I}_{N},$
@@
$\min_{\bm{W}}\operatorname{Tr}\big{[}\bm{W}\hat{\bm{R}}_{x}\bm{W}^{\mathsf{H}}-\bm{W}\hat{\bm{R}}_{xs}-\hat{\bm{R}}^{\mathsf{H}}_{xs}\bm{W}^{\mathsf{H}}+\hat{\bm{R}}_{s}\big{]}+\epsilon_{0}\operatorname{Tr}\big{[}\bm{W}\bm{W}^{\mathsf{H}}\big{]}.$
@@
${\hat{\bm{R}}}_{x}-\epsilon_{0}\bm{F}\preceq\bm{R}_{x}\preceq{\hat{\bm{R}}}_{x}+\epsilon_{0}\bm{F}$
for some $\bm{F}\succeq\bm{0}$.
@@
This is due to Lemma 1 and Theorem 1. Just note that $\operatorname{Tr}\big{[}\bm{W}(\hat{\bm{R}}_{x}+\epsilon_{0}\bm{F})\bm{W}^{\mathsf{H}}-\bm{W}\hat{\bm{R}}_{xs}-\hat{\bm{R}}^{\mathsf{H}}_{xs}\bm{W}^{\mathsf{H}}+\hat{\bm{R}}_{s}\big{]}=\\ \operatorname{Tr}\big{[}\bm{W}\hat{\bm{R}}_{x}\bm{W}^{\mathsf{H}}-\bm{W}\hat{\bm{R}}_{xs}-\hat{\bm{R}}^{\mathsf{H}}_{xs}\bm{W}^{\mathsf{H}}+\hat{\bm{R}}_{s}\big{]}+\epsilon_{0}\operatorname{Tr}\big{[}\bm{W}\bm{F}\bm{W}^{\mathsf{H}}\big{]}$. This completes the proof. $\square$
Note that in Theorem 2, the second-order moment of $\mathbf{s}$ is not considered because it does not influence the optimal solution of $\bm{W}$: i.e., the optimal solution of $\bm{W}$ does not depend on the value of $\bm{R}_{s}$. Theorem 2 gives a new theoretical interpretation of the popular ridge regression in machine learning from the perspective of distributional robustness against second-moment uncertainties of the feature vector $\mathbf{x}$; another interpretation of ridge regression from the perspective of distributional robustness under martingale constraints is identified in [31, Ex. 3.3]. When the uncertainty is quantified by the Wasserstein distance, a similar result can be seen in [32, Prop. 3], [33, Prop. 2], which however is not a ridge regression formulation because in [32, Prop. 3] and [33, Prop. 2], the loss function is square-rooted and the norm regularizer is not squared; cf. also [27, Rem. 18 and 19]. The corollary below justifies the rationale of any norm-regularized method.
@@
Motivated by Theorem 2, the following corollary is immediate, which gives another interpretation of ridge regression and Tikhonov regularization from the perspective of data augmentation through data perturbation (cf. noise injection in image [34] and speech [35] processing).
@@
Suppose that $\mathbf{\Delta}_{x}$ is uncorrelated with $\mathbf{x}$, with $\mathbf{s}$, and with $\mathbf{\Delta}_{s}$; in addition, $\mathbf{\Delta}_{s}$ is uncorrelated with $\mathbf{x}$. If the second-order moment of $\mathbf{\Delta}_{x}$ is upper bounded as $\mathbb{E}\mathbf{\Delta}_{x}\mathbf{\Delta}^{\mathsf{H}}_{x}\preceq\epsilon_{0}\bm{I}_{N}$, then the distributionally robust linear estimator of $\bm{W}$ becomes a ridge regression (i.e., squared-$F$-norm regularized) method
$\min_{\bm{W}}\operatorname{Tr}\big{[}\bm{W}\hat{\bm{R}}_{x}\bm{W}^{\mathsf{H}}-\bm{W}\hat{\bm{R}}_{xs}-\hat{\bm{R}}^{\mathsf{H}}_{xs}\bm{W}^{\mathsf{H}}+\hat{\bm{R}}_{s}\big{]}+\epsilon_{0}\operatorname{Tr}\big{[}\bm{W}\bm{W}^{\mathsf{H}}\big{]}.$
The regularization term becomes $\operatorname{Tr}\big{[}\bm{W}\bm{F}\bm{W}^{\mathsf{H}}\big{]}$, which is known as the Tikhonov regularizer, if $\mathbb{E}\mathbf{\Delta}_{x}\mathbf{\Delta}^{\mathsf{H}}_{x}\preceq\epsilon_{0}\bm{F}$, for some $\bm{F}\succeq\bm{0}$. $\square$
@@
IV-B Complex Uncertainty Sets
Below we remark on more general construction methods for the uncertainty set of $\bm{R}$ using the Wasserstein distance and the $F$-norm, beyond the moment-based methods in Definitions 1$\sim$6. However, note that such complicated approaches are computationally prohibitive in practice when $N$ or $M$ is large.
@@ IV-B1 Wasserstein Distributionally Robust Beamforming
We start with the Wasserstein distance:
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}}&\operatorname{Tr}\big{[}-\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{-1}_{x}\bm{R}_{xs}+\bm{R}_{s}\big{]}\\ \text{s.t.}&\operatorname{Tr}\left[{\bm{R}+\hat{\bm{R}}-2(\hat{\bm{R}}^{1/2}\bm{R}\hat{\bm{R}}^{1/2})^{1/2}}\right]\leq\epsilon^{2}_{0}\\ &\bm{R}\succeq\bm{0},~{}\bm{R}_{x}\succ\bm{0}.\end{array}$ (46)
The first constraint in the above display is a particularization of the Wasserstein distance between $\mathcal{CN}(\bm{0},\bm{R})$ and $\mathcal{CN}(\bm{0},\hat{\bm{R}})$. Problem (46) is equivalent, in the sense of the same maximizer(s), to
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}}&\operatorname{Tr}\bm{R}\\ \text{s.t.}&\operatorname{Tr}\left[{\bm{R}+\hat{\bm{R}}-2(\hat{\bm{R}}^{1/2}\bm{R}\hat{\bm{R}}^{1/2})^{1/2}}\right]\leq\epsilon^{2}_{0}\\ &\bm{R}\succeq\bm{0},~{}\bm{R}_{x}\succ\bm{0},\end{array}$ (47)
since the objective of (46) is increasing in $\bm{R}$; cf. Theorem 1.
@@
Problem (47) can be equivalently reformulated into a linear P-SDP
$\begin{array}[]{cl}\displaystyle\max_{\bm{R},\bm{U}}&\operatorname{Tr}\bm{R}\\ \text{s.t.}&\operatorname{Tr}\left[{\bm{R}+\hat{\bm{R}}-2\bm{U}}\right]\leq\epsilon^{2}_{0}\\ &\left[\begin{array}[]{cc}\hat{\bm{R}}^{1/2}\bm{R}\hat{\bm{R}}^{1/2}&\bm{U}\\ \bm{U}&\bm{I}_{N+M}\end{array}\right]\succeq\bm{0}\\ &\bm{R}\succeq\bm{0},~{}\bm{R}_{x}\succ\bm{0},~{}\bm{U}\succeq\bm{0}.\end{array}$ (48)
@@
This is by applying the Schur complement. $\square$
Complex-valued linear P-SDP can be solved using, e.g., the YALMIP solver. [^2] See https://yalmip.github.io/inside/complexproblems/.
Suppose that $\bm{R}^{\star}$ solves (48). The corresponding Wasserstein distributionally robust beamformer is given as
$\bm{W}^{\star}_{\text{DR-Wasserstein}}=\bm{R}^{\star\mathsf{H}}_{xs}\bm{R}^{\star-1}_{x}.$ (49)
Next, we separately investigate the uncertainties in $\hat{\bm{R}}_{s}$ and $\hat{\bm{R}}_{v}$. From (36), we have
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}_{s},\bm{R}_{v}}&\operatorname{Tr}\big{[}\bm{R}_{s}-\bm{R}_{s}\bm{H}^{\mathsf{H}}(\bm{H}\bm{R}_{s}\bm{H}^{\mathsf{H}}+\bm{R}_{v})^{-1}\bm{H}\bm{R}_{s}\big{]}\\ \text{s.t.}&\operatorname{Tr}\left[{\bm{R}_{s}+\hat{\bm{R}}_{s}-2(\hat{\bm{R}}_{s}^{1/2}\bm{R}_{s}\hat{\bm{R}}_{s}^{1/2})^{1/2}}\right]\leq\epsilon^{2}_{1}\\ &\operatorname{Tr}\left[{\bm{R}_{v}+\hat{\bm{R}}_{v}-2(\hat{\bm{R}}_{v}^{1/2}\bm{R}_{v}\hat{\bm{R}}_{v}^{1/2})^{1/2}}\right]\leq\epsilon^{2}_{2}\\ &\bm{R}_{s}\succeq\bm{0},~{}\bm{R}_{v}\succeq\bm{0},\end{array}$ (50)
where we ignore the uncertainty of $\bm{H}$ for technical tractability. Problem (50) can be transformed into a linear P-SDP using a similar technique as in Proposition 1. One can just introduce an inequality $\bm{U}\succeq\bm{R}_{s}\bm{H}^{\mathsf{H}}(\bm{H}\bm{R}_{s}\bm{H}^{\mathsf{H}}+\bm{R})^{-1}\bm{H}\bm{R}_{s}$ and the objective function will become $\operatorname{Tr}\left[{\bm{R}_{s}-\bm{U}}\right]$.
Suppose that $(\bm{R}_{s}^{\star},\bm{R}_{v}^{\star})$ solves (50). The corresponding Wasserstein distributionally robust beamformer is given as
$\bm{W}^{\star}_{\text{DR-Wasserstein-Individual}}=\bm{R}_{s}^{\star}\bm{H}^{\mathsf{H}}[\bm{H}\bm{R}_{s}^{\star}\bm{H}^{\mathsf{H}}+\bm{R}_{v}^{\star}]^{-1}.$ (51)